OpenStudy (anonymous):

Which is equal to 83 • 88? A. 85 B. 88 C. 118 D. 811

OpenStudy (anonymous):

whats the largest number?

OpenStudy (anonymous):

7304, none of the answers

OpenStudy (anonymous):

the dot means to multiply

OpenStudy (anonymous):

Yep @sop002 agreed

OpenStudy (anonymous):

None

OpenStudy (anonymous):

ohhh sry it is 8^3 X 8^8

OpenStudy (anonymous):

i say A

OpenStudy (anonymous):

because 8^3 X 8^8=8589934592

OpenStudy (anonymous):

yup thats what i just got so u just took the first two numbers right 8^5?

OpenStudy (anonymous):

lol it have 8^5 and 8^11 here so i dont know whitch one

OpenStudy (anonymous):

which *

OpenStudy (anonymous):

D X^3*X^4=X^7

OpenStudy (anonymous):

when combing terms you add exponents

OpenStudy (anonymous):

@RyleeC rightt

OpenStudy (anonymous):

oh yea 8 plus 3 will be 11

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

smart lol